Re: beam today

Date view Thread view Subject view Author view Attachment view

From: W.J. Llope (llope_at_physics.rice.edu)
Date: Sun Mar 09 2003 - 20:41:31 EST


On Sun, 9 Mar 2003, Xu, Zhangbu wrote:

> Hi,
> We will take TOF data today once the beam is stable.
> I am in the control room.
> In the online display, have we zeroed out the bad channels?

bad channels at this point lead to zero's in the hit patterns...
so no need to zero out anything... the only 'hot' channel i know about
was the (intermittant) tofr ch 38 which i pointed out a few days
ago... also some tofp adcs have high pedestals but these are
not in the range you're looking at...

i just checked - the updated online histograms aren't looking correct
to me - there must be some confusion btw sergey and me and/or a bug...
basically what happened is that i gave the code to sergey on last
wednesday but it sat around until today since he was hosting a workshop.
sergey and i did not get a chance to discuss/iterate and
this is undoubtedly why these plots don't look right just yet.
as you mentioned the online data is o.k. - as expected - so i wouldn't
worry too much about these online plots. i will try to solve
this with sergey tomorrow. i see lijuan has been looking
at the localmon data and that's much more informative anyway...

good luck tonight!
        bill

> Thanks!
> Zhangbu
>

-- 
_________________________________________________________
W.J. Llope, Ph.D.                  Res. Assoc. Professor
                 http://wjllope.rice.edu/default.html
                 llope_at_physics.rice.edu
T.W. Bonner Nuclear Lab.     
Rice University, MS-315      
6100 S. Main                       phone: 713-348-4741  
Houston, TX 77005-1892             fax:   713-348-5215

Date view Thread view Subject view Author view Attachment view

This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:37 EDT